OpenStudy (anonymous):

Marlye posted a story on her blog. On the first day there were 80 responses and on the second day there were 180 responses. What percent is the increase, 100, of the original number of responses, 80? a. 55% b. 80% c. 100% d.125%

OpenStudy (popitree):

change is always calculated based on the initial value so increase in absolute value = 180-80 = 100 so chnage in percentage = absolute change/initial value = 100/80 = 1.25 so in percentage it is 125%